OpenStudy (anonymous):

Can some one help me out?? What are the solutions of the quadratic euation x^2 + 11x= -24 -3,-8 -3,8 3,-8 3,8

OpenStudy (anonymous):

i think b

OpenStudy (unklerhaukus):

you could rearrange it, \[ax^2+bx+c=0\] then apply the quadratic formula\[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

ok thanks

OpenStudy (callisto):

If you are good at factorization, then you can factor it to solve the equation. x^2 + 11x= -24 x^2 + 11x + 24 = 0 Factorize it in the form (x-a)(x-b) = 0 So, either (x-a) = 0 or (x-b) = 0

OpenStudy (anonymous):

A

OpenStudy (anonymous):

No